# Independence Constrained Disentangled Representation Learning from Epistemological Perspective

### v1

@@ 3.3.1 MNIST
We trained the models with 1 ten-dimensional discrete variable, 2 continuous variables, and 62 noise variables. The ideal outcome of Disentangled Representation Learning is that the data encoder will map the discrete variable to the digit class, and the two continuous variables will correspond to the width and rotation of the digit. The results of our experiment are provided in Figure 1. In each row, we keep the discrete variable digit invariant, and traverse on the continuous variable rotation. When the model is trained without the independence constraint (Figure 1(a)), we observe that: 1) The variables digit class and rotation are not fully disentangled in the latent space. For example, while we control the digit to be 5 on the fifth row, some 0 and 6 are generated when traversing on rotation; and 2) the variables width and rotation of the digit are not fully disentangled in the latent space. For example, as highlighted in the first row of Figure 1(a), while we only traverse the variable rotation, the width of the digit is also affected. We highlight this trend on the first row, and similar patterns can be observed on other rows. In contrast, with the enhancement of the independent constraint (Figure 1(b)), both digit and width remain unchanged when traversing on rotation.
@@
Figure 5: FashionMNIST Samples for comparison. (a) Without constraining on the independence between variables, item type is affected by traversing on thickness. (b) With the constraint on the independence between variables, item type is unaffected while traversing on thickness. Both models are trained with 1 ten-dimensional discrete variable, 1 continuous variable, and 62 noise variables.
@@ 3.3.3 FashionMNIST
We trained the models with 1 ten-dimensional discrete variable, 1 continuous variable, and 62 noise variables. Ideally, the discrete variable and the continuous variable will correspond to the item class and the thickness of the image. The results of our experiment are provided in Figure 5. A similar conclusion can be drawn as we did on other datasets above. On each row, we keep the discrete variable item class invariant, and traverse on the continuous variable thickness of the image. When the model is trained without the constraint of the independence between latent variables (Figure 5(a)), we observe that the item is affected by variable thickness, as highlighted on the second row and the fourth row. In contrast, the item type remains unaffected when the model is trained with the enhancement of the independent constraint (Figure 5(b)).
@@ 4.1 Disentangled Representation Learning
Disentangled Representation Learning aims to learn a data encoder that can identify true latent variables that are semantically meaningful. [14] suggested that increasing the weight of the KL regularizer in VAE [19] can benefit the quality of disentangled representation learning. [3] proposed that disentanglement quality can be improved by progressively increasing the bottleneck capacity. FactorVAE [17] and $\beta$-TCVAE [7] both penalize the total correlation [34] between latent variables, while FactorVAE uses density-ration trick for total correlation estimation, $\beta$-TCVAE proposed a biased Monte-Carlo estimator to approximate total correlation. Several other methods were also proposed in the paradigm of VAE [20, 11, 18]. However, [23] claimed that unsupervised disentangled representation learning is impossible without inductive biases.
@@ 5 Conclusion
In this paper, we investigated the prior disagreement on the interrelationships between latent variables in Disentangled Representation Learning, and proposed a novel method to improve the quality of disentanglement. First, we build a conceptual bridge between epistemology and disentangled representation learning, thus clarifying what should and should not be independent in the latent space by introducing a two-level latent space framework based on interdisciplinary theories. Then, after clarifying the legitimacy of applying the independence constraint on the problem of Disentangled Representation Learning, we introduce a novel method that applies the mutual information constraint and independence constraint within the Generative Adversarial Network (GAN) framework. Experiments show that our method consistently achieves better disentanglement performance on multiple evaluation metrics, and Qualitative Evaluation results show that our method leads to an improved quality for controllable generation. Besides, our paper introduced a novel perspective to apply causal models to the field of representation learning, which facilitates the development of explainability of deep learning and holds potential for wide-ranging applications that value explainability, transparency and controllability.

### v2

@@ 3.3.1 MNIST
We trained the models with 1 ten-dimensional discrete variable, 2 continuous variables, and 62 noise variables. The ideal outcome of Disentangled Representation Learning is that the data encoder will map the discrete variable to the digit class, and the two continuous variables will correspond to the width and rotation of the digit. The results of our experiment are provided in Figure 1. In each row, we keep the discrete variable digit invariant and traverse on the continuous variable rotation. When the model is trained without the independence constraint (Figure 1(a)), we observe that: 1) The variables digit class and rotation are not fully disentangled in the latent space. For example, while we control the digit to be 5 on the fifth row, some 0 and 6 are generated when traversing on rotation; and 2) the variables width and rotation of the digit are not fully disentangled in the latent space. For example, as highlighted in the first row of Figure 1(a), while we only traverse the variable rotation, the width of the digit is also affected. We highlight this trend on the first row, and similar patterns can be observed on other rows. In contrast, with the enhancement of the independent constraint (Figure 1(b)), both digit and width remain unchanged when traversing on rotation.
@@
Figure 5: FashionMNIST Samples for comparison. Both models are trained with 1 ten-dimensional discrete variable, 1 continuous variable, and 62 noise variables. (a) Without constraining on the independence between variables, item type is affected by traversing on thickness. (b) With the constraint on the independence between variables, item type is unaffected while traversing on thickness.
@@ 3.3.3 FashionMNIST
We trained the models with 1 ten-dimensional discrete variable, 1 continuous variable, and 62 noise variables. Ideally, the discrete variable and the continuous variable will correspond to the item class and the thickness of the image. The results of our experiment are provided in Figure 5. A similar conclusion can be drawn as we did on other datasets above. On each row, we keep the discrete variable item class invariant and traverse on the continuous variable thickness of the image. When the model is trained without the constraint of the independence between latent variables (Figure 5(a)), we observe that the item is affected by variable thickness, as highlighted on the second row and the fourth row. In contrast, the item type remains unaffected when the model is trained with the enhancement of the independent constraint (Figure 5(b)).
@@ 4.1 Disentangled Representation Learning
Disentangled Representation Learning aims to learn a data encoder that can identify true latent variables that are semantically meaningful. [14] suggested that increasing the weight of the KL regularizer in VAE [19] can benefit the quality of disentangled representation learning. [3] proposed that disentanglement quality can be improved by progressively increasing the bottleneck capacity. FactorVAE [17] and $\beta$-TCVAE [7] both penalize the total correlation [34] between latent variables, while FactorVAE uses a density-ratio trick for total correlation estimation, $\beta$-TCVAE proposed a biased Monte-Carlo estimator to approximate total correlation. Several other methods were also proposed in the paradigm of VAE [20, 11, 18]. However, [23] claimed that unsupervised disentangled representation learning is impossible without inductive biases.
@@ 5 Conclusion
In this paper, we investigated the prior disagreement on the interrelationships between latent variables in Disentangled Representation Learning and proposed a novel method to improve the quality of disentanglement. First, we build a conceptual bridge between epistemology and disentangled representation learning, thus clarifying what should and should not be independent in the latent space by introducing a two-level latent space framework based on interdisciplinary theories. Then, after clarifying the legitimacy of applying the independence constraint on the problem of Disentangled Representation Learning, we introduce a novel method that applies the mutual information constraint and independence constraint within the Generative Adversarial Network (GAN) framework. Experiments show that our method consistently achieves better disentanglement performance on multiple evaluation metrics, and Qualitative Evaluation results show that our method leads to an improved quality for controllable generation. Besides, our paper introduced a novel perspective to apply causal models to the field of representation learning, which facilitates the development of explainability of deep learning and holds potential for wide-ranging applications that value explainability, transparency and controllability.